Write the natural numbers from 102 to 113. What fraction of them are prime numbers?

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Natural numbers from 102 to 113 are
102, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113
Total number of natural numbers given = 12
Number of prime numbers = 4 [103, 107, 109, 113]
∴Required Fraction = 4 / 12 = 1 / 3

Thanks